The height of an object thrown in the air is given by h = t2 – 7t + 3, where h is in feet and t is the time in seconds the object has been in motion. At what time (to the nearest tenth) is the object 11 feet in the air?
Could you help to solve this equation?
Thank you.
Cynthia

Answer by stanbon(75887)

The height of an object thrown in the air is given by h = t2 – 7t + 3, where h is in feet and t is the time in seconds the object has been in motion. At what time (to the nearest tenth) is the object 11 feet in the air?
----
h is the height so make it 11 and solve for "t":
t^2-7t+3 = 11
---
t^2 -7t -8 = 0
(t-8)(t+1) = 0
Positive solution:
t = 8 seconds (time at which the object is at a height of 11 ft.
=====================================
Cheers,
Stan H.